OpenStudy (anonymous):

write an equation in point slope form for the line through the given point with the given slope (8,3);m=5

OpenStudy (welshfella):

use the standard equation y - y1 = m(x - x1) where m = slope , and (x1,y1) is a given point on the line

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my options are y+3=5(x-8) y+3=5x-8 y-3=5(x-8) y-3=5(x+8) @welshfella

OpenStudy (welshfella):

well plug in the values and see if you get one of the options y - 3 = ? what goes on right side?

OpenStudy (welshfella):

its got to be 5( x - something) because m = 5 what is last number in the parentheses?

OpenStudy (welshfella):

what is x1? check out my first post

OpenStudy (welshfella):

(x1,y1) = (8,3)

OpenStudy (anonymous):

its the last one

OpenStudy (anonymous):

can i tell you how i got it?

OpenStudy (welshfella):

no - not the last one

OpenStudy (anonymous):

how?

OpenStudy (welshfella):

how did youou get it?

OpenStudy (anonymous):

y-y8=m (x-x8) y-3=5(x-(-8) y-3=5(x+8)

OpenStudy (welshfella):

no - because x1 is not -8

OpenStudy (welshfella):

check the point given in the question

OpenStudy (anonymous):

ok